# Spin-orbit misalignment and residual eccentricity are evidence that neutron star-black hole mergers form through triple star evolution

### v1

@@ 2.1 Evolution of inner binaries
For the vast majority of binaries that evolve to become NSBHs ($\gtrsim 95\,\%$; F. S. Broekgaarden et al., 2021; Z. Xing et al., 2024; C. Liotine et al., 2025), the BH is expected to form first from the more massive primary star ($m_{1}$), while the progenitor of the NS is still on the MS ($m_{2}$). This implies an intermediate evolutionary stage of an inner binary with a BH and an OB-type stellar companion (“BHOB triple”). We thus construct a two-stage binary evolution model for the inner binaries, as follows.
@@ 2. Stage II)
If the orbit is sufficiently close (with semi-major axis $a_{1}\lesssim 10\,\rm AU$), the expanding primary may at some point overfills its Roche lobe and initiates mass transfer (MT) onto the inner binary companion. To assess its stability, we first check whether the donor has an outer convective or radiative envelope based on the criteria from J. Klencki et al. (2020). For convective donors, we assume that the MT is stable for $q>q_{\rm crit;conv}$, where $q=M_{\rm accretor}/M_{\rm donor}$ at the onset of Roche-lobe overflow (RLOF), and set $q_{\rm crit;conv}=0.8$ (M. S. Hjellming & R. F. Webbink, 1987; K. D. Temmink et al., 2023). For radiative donors, we apply two stability requirements. First, for both Stage I) and Stage II), we require that for the MT to remain stable, the final post-interaction orbit must be wider than $20{\rm\,R_{\odot}}$ (based on the separation limit found in J. Klencki et al., 2025). Second, for stellar accretors (i.e., only Stage I)), we require $q>q_{\rm crit;rad}$, where $q_{\rm crit;rad}$ depends on orbital period at RLOF:
@@
where $q({\rm log}P_{\rm RLOF}/{\rm days})$ is a linear fit to $q=0.35$ at ${\rm log}P_{\rm RLOF}/{\rm days}=1$ and $q=0.6$ at ${\rm log}P_{\rm RLOF}/{\rm days}=3$. This criteria approximately reproduces the MT stability from massive radiative donors found in the COMBINE code for our choice of low accretion efficiency $\beta=0.1$ (see Figs. E.6 – E.9 of C. Schürmann et al., 2025), based on a detailed study by C. Schürmann & N. Langer (2024). The origin of MT instability here is a runaway expansion of the mass gainer star when the accretion rate exceeds a certain threshold (see also J. Henneco et al., 2024; M. Y. M. Lau et al., 2024). For a larger accretion efficiency, the mass transfer would be less stable, particularly for periods $P_{\rm RLOF}>100$ days (C. Schürmann et al., 2025). We assume that unstable MT with a radiative donor always leads to a stellar merger (P. Podsiadlowski, 2001; M. U. Kruckow et al., 2016; J. Klencki et al., 2021; P. Marchant et al., 2021). For unstable MT with a convective donor, we consider common envelope (CE) evolution. For Stage I), we apply the alpha-lambda energy formalism (E. P. J. van den Heuvel, 1976; R. F. Webbink, 1984) to determine whether the system survives the CE phase as well as the post-CE orbit. We assume $\alpha_{\rm CE}=1$ and include 100 % contribution from internal and recombination energy into the envelope binding energy calculation (for details see J. Klencki et al., 2021). For Stage II), we assume that all CE events lead to surviving binaries (BH + He-star systems) with a post-CE semi-major axis $a_{\rm post-CE}=4{\rm\,R_{\odot}}$, i.e., sufficiently close to produce a GW-driven NSBH merger even without the aid of the tertiary companion. Our choice to assume 100 % survival rate of CE events for Stage II) allows us to gauge the upper limit for the contribution of the CE channel to NSBH mergers. The choice to fix $a_{\rm post-CE}$ is motivated by the fact that the exact final separation of post-CE systems is poorly known as it is likely determined by the complex interaction between the binary orbit and the circumbinary matter ejected during the dynamical phase of CE (e.g. D. Gagnier & O. Pejcha, 2023; D. Wei et al., 2024). Moreover, in this work they merely serve as a reference to compare our tertiary-driven NSBH mergers to a model where NSBHs could even merge in isolation. We note that a difference choice of $a_{\rm post-CE}$ would result in a somewhat different distribution of BH spin-orbit misalignments from the CE channel: the wider the orbit, the easier it is to tilt by the natal kick associated with NS formation, see Sec. 2.2 (G. Fragione et al., 2021).
For mass transfer events that remain stable, we follow the method detailed in Sec. 2 and App. A of J. Klencki et al. (2025). Briefly, we calculate the change in orbital semi-major axis as:
@@
where $a_{\rm ini}$ ($a_{\rm fin}$) is the semi-major axis before (after) mass transfer and $M_{\rm don}$ ($M_{\rm acc}$) is the mass of the donor (accretor). For the accretion efficiency we set $\beta=0.1$ for Stage I) and $\beta=0.0$ for Stage II). The first choice roughly approximates results from detailed models with rotationally-limited accretion (e.g. N. Langer et al. 2003; M. Renzo & Y. Götberg 2021, although $\beta>0.1$ may be required for some systems, see S. Vinciguerra et al. 2020; T. Lechien et al. 2025), the second is motivated by the Eddington accretion limit. We assume that the non-accreted mass is ejected from the proximity of the accretor ($\gamma=M_{\rm don}/M_{\rm acc}$). For case B/C mass transfer, we assume that the donor transfers 90 % of its envelope ($M_{\rm don;fin}=M_{\rm He;core}+0.1M_{\rm env}$). For case A, this amount is less certain due to the receding convective core (M. Shikauchi et al., 2024) and we calculate $M_{\rm don;fin}$ based on analytical fits to detailed binary evolution models provided by C. Schürmann et al. (2024, their SMC variant). Once a helium star forms as a result of mass transfer, we estimate its further mass loss by applying the wind from W. R. Hamann et al. (1995) lowered by a factor of 10 (S. C. Yoon et al., 2006).
@@
In order to assess the final fate of a star, we apply the birth function for BH and NS formation in close binaries from S. E. Woosley et al. (2020). As their results are based on naked helium star models, we map their pre-supernova masses to our final helium core masses. In addition, following J. R. Hurley et al. (2002), we assume that stars with helium cores between $1.6$ and $2.3{\rm\,M_{\odot}}$ form a NS through an electron-capture supernova (ECSN). We note that P. Podsiadlowski et al. (2004) argued for a wider mass range for ECSN for stars evolving in close binaries, whereas R. Willcox et al. (2021) showed that a narrowed mass range for ECSN from single stars might be required to match the number of low-velocity pulsars. Regardless of the SN type, we assume that all NSs form with a mass of $m_{\rm NS}=1.5{\rm\,M_{\odot}}$. For BH formation, if the collapsing star still retains some of its hydrogen envelope, we assume that 50 % of this envelope is accreted onto the BH and the rest is ejected. Guided by S. E. Woosley et al. (2020), we assume that naked helium stars collapse directly to form BHs with only $5\,\%$ of their mass lost in neutrinos. Besides the mass loss, we assume weak BH natal kicks drawn from a Maxwellian distribution with velocity dispersion $\sigma_{k}=5\,\rm{km\,s^{-1}}$. This is consistent with theoretical predictions for neutrino-driven kicks (H.-T. Janka & D. Kresse, 2024) which struggle to produce kicks $>100\,\rm{km\,s^{-1}}$ even for large neutrino emission asymmetries (A. Burrows et al., 2024), as is supported by the recently discovered BH binary VFTS 243 (A. Vigna-Gómez et al., 2024). The NS natal kicks are expected to be larger and play a significant role in the survival of triples. We therefore consider several model variations for the NS kick magnitudes $v_{k}$:
@@
    Maxwellian_265_30: Maxwellian with $\sigma_{k}=265\rm\,{km\,s^{-1}}$ (G. Hobbs et al., 2005). For ECSN, we instead apply $\sigma_{k}=30\rm\,{km\,s^{-1}}$ (S. Jones et al., 2016; A. Gessner & H.-T. Janka, 2018).
@@
    Igoshev: Maxwellian with $\sigma_{k}=336\rm\,{km\,s^{-1}}$ in 80% of cases and $\sigma_{k}=45\rm\,{km\,s^{-1}}$ 20% otherwise, randomly drawn (A. P. Igoshev et al., 2021), see also (F. Verbunt et al., 2017). No separate ECSN kick treatment.
@@
    Valli: Maxwellian with $\sigma_{k}=100\rm\,{km\,s^{-1}}$ for stars that underwent mass transfer, otherwise $\sigma_{k}=265\rm\,{km\,s^{-1}}$. This is motivated by R. Valli et al. (2025) but is only a rough approximation: in particular we do not include their model for anisotropic kicks (see Sec. 2.2). For ECSN, we instead apply a Maxwellian with $\sigma_{k}=10\rm\,{km\,s^{-1}}$, regardless of mass transfer history. [^1] The assumption of $\sigma_{k}=10\rm\,{km\,s^{-1}}$ to all ECSN may not be a good description of the results from R. Valli et al. (2025).
@@
    Disberg: log-normal distribution with $\mu=5.6$ and $\sigma=0.68$ (P. Disberg & I. Mandel, 2025). No separate ECSN kick treatment.
@@
    Bray: deterministic kick expressed as $v_{k}=\alpha(M_{\rm ejecta}/M_{\rm NS})+\beta$, where $\alpha=115\rm\,{km\,s^{-1}}$ $\beta=10\rm\,{km\,s^{-1}}$, and $M_{\rm ejecta}$ is the mass lost upon the NS formation (J. C. Bray & J. J. Eldridge, 2018; S. M. Richards et al., 2023). No separate ECSN kick treatment.
@@
2.2 Impact of supernovae on triple orbits
@@
where the kick velocity vector $\mathbf{v}_{k}$ points in a isotropically random direction and has a magnitude drawn from either one of the distribution described in Sec. 2.1. Furthermore, the star undergoing SN is subject to sudden mass-loss given by
@@
The mass-loss and natal kick associated with the SN also imparts a kick to the centre-of-mass velocity of the inner binary which modifies the orbit of the outer binary and potentially disrupts it. We update the outer orbit in an analogue way to the inner binary, where the new relative velocity vector of the outer binary and total mass of the inner binary are given by (C. X. Lu & S. Naoz, 2019) [^2] Different to Eq. (24) of C. X. Lu & S. Naoz (2019) we find a negative sign in front of the $\mathbf{v}_{k}$-term in Eq. (5), which is irrelevant for isotropically random kick directions. {widetext}
@@
In Sec. 3, we show that for the NS formation these constrains critically shape the distributions of surviving triples with an NSBH in the inner binary and their long-term dynamical evolution towards becoming a GW source.
2.3 Gravitational three-body dynamics
For any triple that successfully formed an NSBH in the inner binary without being disrupted or becoming dynamically unstable, we integrate the long-term secular three-body dynamics as follows. From the post-SN orbital parameters we construct the eccentricity vectors $\mathbf{e}_{1(2)}$ and dimensionless orbital angular momentum vectors $\mathbf{j}_{1(2)}$ of the inner and outer orbit, respectively (see Fig. 1). The eccentricity vectors are defined to point towards the orbital periapsis and have lengths $|\mathbf{e}_{1(2)}|={e}_{1(2)}$. The dimensionless orbital angular momentum vectors $\mathbf{j}_{1(2)}$ are perpendicular to the orbital plane of the inner and outer binary, respectively, and have lengths $|\mathbf{j}_{1(2)}|=\sqrt{1-{e}_{1(2)}^{2}}$. Additionally, we consider the spin vector of the BH $\mathbf{S}_{\rm BH}=\mathbf{\chi}_{\rm BH}Gm_{\rm BH}^{2}/c$, where we assume a dimensionless spin parameter $\mathbf{\chi}_{\rm BH}=0.25$ consistent with the low spin magnitudes inferred from previous BH merger events (R. Abbott et al., 2023) and we assume that $\mathbf{S}_{\rm BH}$ is initially aligned with the pre-SN inner orbital angular momentum, as a result of stellar evolution. We consider the equations of motion for $\mathbf{e}_{1(2)}$, $\mathbf{j}_{1(2)}$, $\mathbf{S}_{\rm BH}$, and $a_{1}$ given by the ZKL effect to the octupole order accounting for the Newtonian dynamics of hierarchical three-body systems, GW radiation and Schwarzschild precession of the inner binary, and relativistic spin-orbit coupling. That is, we include Eqs. (17) – (20) of B. Liu et al. (2015) and Eqs. (13), (15) – (17), and (25) – (26) of C. L. Rodriguez & F. Antonini (2018), which are lengthy and not repeated here for convenience. We integrate them over time using the Radau method (E. Hairer & G. Wanner, 1996) with relative and absolute tolerances of $10^{-10}$. The integrations are terminated if either the NSBHs merge or if the integration time exceeds $\min(10^{3}\times t_{\rm ZKL},13\,\rm Gyr)$, where the ZKL timescale of the leading-order quadrupole effect is given by

### v2

@@ II.1 Evolution of inner binaries
For the vast majority of binaries that evolve to become NSBHs ($\gtrsim 95\,\%$; Broekgaarden et al., 2021; Xing et al., 2024; Liotine et al., 2025), the BH is expected to form first from the more massive primary star ($m_{1}$), while the progenitor of the NS is still on the MS ($m_{2}$). This implies an intermediate evolutionary stage of an inner binary with a BH and an OB-type stellar companion (“BHOB triple”). We thus construct a two-stage binary evolution model for the inner binaries, as follows.
@@ 2. Stage II)
If the orbit is sufficiently close (with semi-major axis $a_{1}\lesssim 10\,\rm AU$), the expanding primary may at some point overfills its Roche lobe and initiates mass transfer (MT) onto the inner binary companion. To assess its stability, we first check whether the donor has an outer convective or radiative envelope based on the criteria from Klencki et al. (2020). For convective donors, we assume that the MT is stable for $q>q_{\rm crit;conv}$, where $q=M_{\rm accretor}/M_{\rm donor}$ at the onset of Roche-lobe overflow (RLOF), and set $q_{\rm crit;conv}=0.8$ (Hjellming and Webbink, 1987; Temmink et al., 2023). For radiative donors, we apply two stability requirements. First, for both Stage I) and Stage II), we require that for the MT to remain stable, the final post-interaction orbit must be wider than $20{\rm\,R_{\odot}}$ (based on the separation limit found in Klencki et al., 2025). Second, for stellar accretors (i.e., only Stage I)), we require $q>q_{\rm crit;rad}$, where $q_{\rm crit;rad}$ depends on orbital period at RLOF:
@@
where $q({\rm log}P_{\rm RLOF}/{\rm days})$ is a linear fit to $q=0.35$ at ${\rm log}P_{\rm RLOF}/{\rm days}=1$ and $q=0.6$ at ${\rm log}P_{\rm RLOF}/{\rm days}=3$. This criteria approximately reproduces the MT stability from massive radiative donors found in the COMBINE code for our choice of low accretion efficiency $\beta=0.1$ (see Figs. E.6 – E.9 of Schürmann et al., 2025), based on a detailed study by Schürmann and Langer (2024). The origin of MT instability here is a runaway expansion of the mass gainer star when the accretion rate exceeds a certain threshold (see also Henneco et al., 2024; Lau et al., 2024). For a larger accretion efficiency, the mass transfer would be less stable, particularly for periods $P_{\rm RLOF}>100$ days (Schürmann et al., 2025). We assume that unstable MT with a radiative donor always leads to a stellar merger (Podsiadlowski, 2001; Kruckow et al., 2016; Klencki et al., 2021; Marchant et al., 2021). For unstable MT with a convective donor, we consider common envelope (CE) evolution. For Stage I), we apply the alpha-lambda energy formalism (van den Heuvel, 1976; Webbink, 1984) to determine whether the system survives the CE phase as well as the post-CE orbit. We assume $\alpha_{\rm CE}=1$ and include 100 % contribution from internal and recombination energy into the envelope binding energy calculation (for details see Klencki et al., 2021). For Stage II), we assume that all CE events lead to surviving binaries (BH + He-star systems) with a post-CE semi-major axis $a_{\rm post-CE}=4{\rm\,R_{\odot}}$, i.e., sufficiently close to produce a GW-driven NSBH merger even without the aid of the tertiary companion. Our choice to assume 100 % survival rate of CE events for Stage II) allows us to gauge the upper limit for the contribution of the CE channel to NSBH mergers. The choice to fix $a_{\rm post-CE}$ is motivated by the fact that the exact final separation of post-CE systems is poorly known as it is likely determined by the complex interaction between the binary orbit and the circumbinary matter ejected during the dynamical phase of CE (e.g. Gagnier and Pejcha, 2023; Wei et al., 2024). Moreover, in this work they merely serve as a reference to compare our tertiary-driven NSBH mergers to a model where NSBHs could even merge in isolation. We note that a difference choice of $a_{\rm post-CE}$ would result in a somewhat different distribution of BH spin-orbit misalignments from the CE channel: the wider the orbit, the easier it is to tilt by the natal kick associated with NS formation, see Sec. II.2 (Fragione et al., 2021).
For mass transfer events that remain stable, we follow the method detailed in Sec. 2 and App. A of Klencki et al. (2025). Briefly, we calculate the change in orbital semi-major axis as:
@@
where $a_{\rm ini}$ ($a_{\rm fin}$) is the semi-major axis before (after) mass transfer and $M_{\rm don}$ ($M_{\rm acc}$) is the mass of the donor (accretor). For the accretion efficiency we set $\beta=0.1$ for Stage I) and $\beta=0.0$ for Stage II). The first choice roughly approximates results from detailed models with rotationally-limited accretion (e.g. Langer et al. 2003; Renzo and Götberg 2021, although $\beta>0.1$ may be required for some systems, see Vinciguerra et al. 2020; Lechien et al. 2025), the second is motivated by the Eddington accretion limit. We assume that the non-accreted mass is ejected from the proximity of the accretor ($\gamma=M_{\rm don}/M_{\rm acc}$). For case B/C mass transfer, we assume that the donor transfers 90 % of its envelope ($M_{\rm don;fin}=M_{\rm He;core}+0.1M_{\rm env}$). For case A, this amount is less certain due to the receding convective core (Shikauchi et al., 2024) and we calculate $M_{\rm don;fin}$ based on analytical fits to detailed binary evolution models provided by Schürmann et al. (2024, their SMC variant). Once a helium star forms as a result of mass transfer, we estimate its further mass loss by applying the wind from Hamann et al. (1995) lowered by a factor of 10 (Yoon et al., 2006).
@@
In order to assess the final fate of a star, we apply the birth function for BH and NS formation in close binaries from Woosley et al. (2020). As their results are based on naked helium star models, we map their pre-supernova masses to our final helium core masses. In addition, following Hurley et al. (2002), we assume that stars with helium cores between $1.6$ and $2.3{\rm\,M_{\odot}}$ form a NS through an electron-capture supernova (ECSN). We note that Podsiadlowski et al. (2004) argued for a wider mass range for ECSN for stars evolving in close binaries, whereas Willcox et al. (2021) showed that a narrowed mass range for ECSN from single stars might be required to match the number of low-velocity pulsars. Regardless of the SN type, we assume that all NSs form with a mass of $m_{\rm NS}=1.5{\rm\,M_{\odot}}$. For BH formation, if the collapsing star still retains some of its hydrogen envelope, we assume that 50 % of this envelope is accreted onto the BH and the rest is ejected. Guided by Woosley et al. (2020), we assume that naked helium stars collapse directly to form BHs with only $5\,\%$ of their mass lost in neutrinos. Besides the mass loss, we assume weak BH natal kicks drawn from a Maxwellian distribution with velocity dispersion $\sigma_{k}=5\,\rm{km\,s^{-1}}$. This is consistent with theoretical predictions for neutrino-driven kicks (Janka and Kresse, 2024) which struggle to produce kicks $100\,\rm{km\,s^{-1}}$ even for large neutrino emission asymmetries (Burrows et al., 2024), as is supported by the recently discovered BH binary VFTS 243 (Vigna-Gómez et al., 2024). The NS natal kicks are expected to be larger and play a significant role in the survival of triples. We therefore consider several model variations for the NS kick magnitudes $v_{k}$:
@@
    Maxwellian_265_30: Maxwellian with $\sigma_{k}=265\rm\,{km\,s^{-1}}$ (Hobbs et al., 2005). For ECSN, we instead apply $\sigma_{k}=30\rm\,{km\,s^{-1}}$ (Jones et al., 2016; Gessner and Janka, 2018).
@@
    Igoshev: Maxwellian with $\sigma_{k}=336\rm\,{km\,s^{-1}}$ in 80% of cases and $\sigma_{k}=45\rm\,{km\,s^{-1}}$ 20% otherwise, randomly drawn (Igoshev et al., 2021), see also (Verbunt et al., 2017). No separate ECSN kick treatment.
@@
    Valli: Maxwellian with $\sigma_{k}=100\rm\,{km\,s^{-1}}$ for stars that underwent mass transfer, otherwise $\sigma_{k}=265\rm\,{km\,s^{-1}}$. This is motivated by Valli et al. (2025) but is only a rough approximation: in particular we do not include their model for anisotropic kicks (see Sec. II.2). For ECSN, we instead apply a Maxwellian with $\sigma_{k}=10\rm\,{km\,s^{-1}}$, regardless of mass transfer history. [^1] The assumption of $\sigma_{k}=10\rm\,{km\,s^{-1}}$ to all ECSN may not be a good description of the results from Valli et al. (2025).
@@
    Disberg: log-normal distribution with $\mu=5.6$ and $\sigma=0.68$ (Disberg and Mandel, 2025). No separate ECSN kick treatment.
@@
    Bray: deterministic kick expressed as $v_{k}=\alpha(M_{\rm ejecta}/M_{\rm NS})+\beta$, where $\alpha=115\rm\,{km\,s^{-1}}$ $\beta=10\rm\,{km\,s^{-1}}$, and $M_{\rm ejecta}$ is the mass lost upon the NS formation (Bray and Eldridge, 2018; Richards et al., 2023). No separate ECSN kick treatment.
@@
II.2 Impact of supernovae on triple orbits
@@
where the kick velocity vector $\mathbf{v}_{k}$ points in a isotropically random direction and has a magnitude drawn from either one of the distribution described in Sec. II.1. Furthermore, the star undergoing SN is subject to sudden mass-loss given by
@@
The mass-loss and natal kick associated with the SN also imparts a kick to the centre-of-mass velocity of the inner binary which modifies the orbit of the outer binary and potentially disrupts it. We update the outer orbit in an analogue way to the inner binary, where the new relative velocity vector of the outer binary and total mass of the inner binary are given by (Lu and Naoz, 2019) [^2] Different to Eq. (24) of Lu and Naoz (2019) we find a negative sign in front of the $\mathbf{v}_{k}$-term in Eq. (5), which is irrelevant for isotropically random kick directions.
@@
In Sec. III, we show that for the NS formation these constrains critically shape the distributions of surviving triples with an NSBH in the inner binary and their long-term dynamical evolution towards becoming a GW source.
II.3 Gravitational three-body dynamics
For any triple that successfully formed an NSBH in the inner binary without being disrupted or becoming dynamically unstable, we integrate the long-term secular three-body dynamics as follows. From the post-SN orbital parameters we construct the eccentricity vectors $\mathbf{e}_{1(2)}$ and dimensionless orbital angular momentum vectors $\mathbf{j}_{1(2)}$ of the inner and outer orbit, respectively (see Fig. 1). The eccentricity vectors are defined to point towards the orbital periapsis and have lengths $|\mathbf{e}_{1(2)}|={e}_{1(2)}$. The dimensionless orbital angular momentum vectors $\mathbf{j}_{1(2)}$ are perpendicular to the orbital plane of the inner and outer binary, respectively, and have lengths $|\mathbf{j}_{1(2)}|=\sqrt{1-{e}_{1(2)}^{2}}$. Additionally, we consider the spin vector of the BH $\mathbf{S}_{\rm BH}=\mathbf{\chi}_{\rm BH}Gm_{\rm BH}^{2}/c$, where we assume a dimensionless spin parameter $\mathbf{\chi}_{\rm BH}=0.25$ consistent with the low spin magnitudes inferred from previous BH merger events (Abbott et al., 2023) and we assume that $\mathbf{S}_{\rm BH}$ is initially aligned with the pre-SN inner orbital angular momentum, as a result of stellar evolution. We consider the equations of motion for $\mathbf{e}_{1(2)}$, $\mathbf{j}_{1(2)}$, $\mathbf{S}_{\rm BH}$, and $a_{1}$ given by the ZKL effect to the octupole order accounting for the Newtonian dynamics of hierarchical three-body systems, GW radiation and Schwarzschild precession of the inner binary, and relativistic spin-orbit coupling. That is, we include Eqs. (17) – (20) of Liu et al. (2015) and Eqs. (13), (15) – (17), and (25) – (26) of Rodriguez and Antonini (2018), which are lengthy and not repeated here for convenience. We integrate them over time using the Radau method (Hairer and Wanner, 1996) with relative and absolute tolerances of $10^{-10}$. The integrations are terminated if either the NSBHs merge or if the integration time exceeds $\min(10^{3}\times t_{\rm ZKL},13\,\rm Gyr)$, where the ZKL timescale of the leading-order quadrupole effect is given by